OpenStudy (anonymous):

PLEASE HELP? I dont think anyone can get this i asked and asked!!! Please try best answer?fan?asap Describe an infinite geometric series with a beginning value of 2 that converges to 10. What are the first 4 terms of the series?

OpenStudy (phi):

See the formula for summing the geometric series http://en.wikipedia.org/wiki/Geometric_series#Formula when n= infinity and r is less than 1, r^n approaches zero the sum of an infinite number of terms becomes \[ S= \sum_{n=0}^{\infty} ar^n= \frac{a}{1-r} \]

OpenStudy (phi):

a is the first term. they tell you that. S is the sum of an infinite number of terms. they tell you that. now use algebra to solve for r.

OpenStudy (anonymous):

thats what i dont get it why is there "2 converge to 10" thats what confuses me you dont have to give me the answer but a little more hints would be greatly apprieciated

OpenStudy (phi):

You have to know about a geometric series For example, there is this series It starts with 1/2 as the first term \[ \frac{1}{2} \] then we set the next term to 1/2 times the first term. 1/2 * 1/2 is 1/4 \[ \frac{1}{2} + \frac{1}{4}\] the 3rd term is 1/2 times the second term. 1/2 * 1/4 = 1/8 \[ \frac{1}{2} + \frac{1}{4} + \frac{1}{8} + ...\] we do that forever

OpenStudy (phi):

Describe an infinite geometric series with a beginning value of 2 that converges to 10 that is asking for a series whose first term is 2 2 and we multiply the 2 times a factor (which we don't know YET, but will call r) the first two terms are: \[ 2 + 2r \] the 3rd term is r times the 2nd term: r * 2r = 2r^2 \[ 2 + 2r + 2r^2 + ....\] we keep doing that forever but when we add them all up, they add up to 10 (the more terms we add the closer we get to 10, but we never *quite* get there) the series converges to 10

OpenStudy (anonymous):

meaning i only plug in 2? or 2-10? im sorry i am ust confused about the question but i actually understand evrything your saying just not what the question wants me to do

OpenStudy (phi):

luckily we have a formula for what the series converges (adds up) to: \[ 2 + 2r + 2r^2 + 2r^3 + 2r^4 + ....= \frac{2}{1-r} \] they tell us they want this thing to add up to 10 we can use the formula: \[ \frac{2}{1-r} = 10 \] I would "flip" both sides (remember 10 can be thought of as 10/1) can you do that ?

OpenStudy (phi):

you want to solve for "r" there are a few ways to do this.

OpenStudy (phi):

are you stuck on the algebra ?

OpenStudy (anonymous):

would it be r=8?

OpenStudy (anonymous):

i multiplied 1-r to 10 which led me to ------ 2= 10 * 1-r.... then i got 2=10-r. then, -8=-r then divided by -1

OpenStudy (phi):

no, but you can check your answer this way: replace r with 8 in the equation \[\frac{2}{1-8} = ?10 \] 1-8 is -7 and 2/-7 is not 10

OpenStudy (anonymous):

ill try again

OpenStudy (phi):

oh, ok, when you multiply 1-r it is one "chunk" so put it in parens you get 2 = 10*(1-r) you have to distribute the 10. (multiply BOTH 1 and -r by 10) But I would not bother I would first, divide both sides by 10

OpenStudy (anonymous):

ill tell you what i just got but it looks backwards.... r= -19

OpenStudy (phi):

write down your steps, and I'll tell you if you are on the right track

OpenStudy (anonymous):

this time i multiplied 10 by 2 nd got ... 1-r= 10*2, then, 1-r=20, subtract 1 and get .....-r=19, then divide and get r= -19

OpenStudy (phi):

pay attention. we start at the beginning \[ \frac{2}{1-r} = 10 \] multiply both sides by (1-r) \[ \frac{2}{\cancel{1-r}} \cdot \frac{\cancel{(1-r)}}{1} = 10 \cdot (1-r)\]

OpenStudy (phi):

now divide both sides by 10. can you do that ?

OpenStudy (phi):

like this \[\frac{2}{10} = \cancel{\frac{10}{10}} \cdot (1-r) \]

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (phi):

of course 2/10 is 1/5 \[ \frac{1}{5}= 1-r \]

OpenStudy (anonymous):

so i could subtract 1 from1/5 next? to get it isolated?

OpenStudy (phi):

yes, we could add -1 (or subtract 1, but I always think "ADD" and use negative numbers) -1+1/5 = 1-1 -r

OpenStudy (phi):

we get -1 + 1/5 = -r it looks nicer if we multiply both sides (ALL TERMS) by -1 -1 * (-1+1/5) = -1*-r +1 -1/5 = r

OpenStudy (anonymous):

4/5?

OpenStudy (phi):

yes, 4/5 you can check 2/(1-4/5) is 2/(1/5) which is 2*5 = 10 now can you write down the first few terms? the first one is 2 2 then 2nd term is 2 * 4/5 the third term is the 2nd term times 4/5 and so on If you are ambitious, you can write down all of them, add them up and see if you get 10 (haha)

OpenStudy (phi):

btw, try some of Khan's videos http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/basic-equation-practice/v/equations-3 because solving equations is important.

OpenStudy (anonymous):

@phi thank you so much your the best and i loved how you helped and explain really thank you!!

OpenStudy (anonymous):

can i ask one more question? i want to see if i got it right

OpenStudy (phi):

If you have time, start with this one http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/equations_beginner/v/simple-equations and watch the next few.

OpenStudy (phi):

?

OpenStudy (anonymous):

Justine earned $26,000 during the first year of her job at city hall. After each year she received a 3% raise. Find her total earnings during the first five years on the job. $138,037.53***** $1,004,704.20 $4,020.51 $108,774.30 am i right?

OpenStudy (anonymous):

and i sure will!!

OpenStudy (phi):

this is a geometric series (what a coincidence). the starting number is 26000 you multiply by 1.03 ( 0.03 is the raise 1.03 is the salary plus raise) if we write it out: 26000 + 26000*1.03 + 26000*1.03^2 + 26000*1.03^3+ 26000*1.03^4 we could just add it up, but it's better to use the formula \[ \sum_{n=0}^{4} 26000 (1.03)^n = 26000 \cdot \frac{1-(1.03)^5}{1-(1.03)}\] the bottom is 1-1.03 = -.03 the top is 1 - 1.159274 = -0.159274 the sum is \[ 26000 \cdot \frac{ -0.159274}{-0.03} = 26000*5.309135 = 138037.53

OpenStudy (phi):

the sum is \[ 26000 \cdot \frac{ -0.159274}{-0.03} = 26000*5.309135 = 138037.53 \]

OpenStudy (anonymous):

so i was right?

OpenStudy (phi):

I got 138037.53 is that what you picked ?

OpenStudy (anonymous):

yeah thats what i picked thats why i put the star and you were right thank you soooooo much:)